OpenStudy (helpblahblahblah):

ILL GIVE MEDALS: 20 is 4% of what? A.50 B. 200 C. 500 D. 800

OpenStudy (helpblahblahblah):

i can figure it out please explain

OpenStudy (helpblahblahblah):

can you help @beastmachine

OpenStudy (anonymous):

one sec

OpenStudy (helpblahblahblah):

k

OpenStudy (destinymasha):

i know its 500, but how do you find out?

OpenStudy (destinymasha):

http://www.mathsisfun.com/percentage.html

OpenStudy (anonymous):

there is a formula for this... is/of = %/100 20 is 4% of what.. 20/x = 4/100 cross multiply 4x = 100(20) 4x = 2000 x = 2000/4 x = 500 Understand how I set it up ?

OpenStudy (helpblahblahblah):

yeah thanks

OpenStudy (anonymous):

no problem :)